The skydriver is falling with a force of 2500 N and an upwards force of 450 N. Calculate the net force
1
Expert's answer
2021-09-28T11:13:41-0400

Given:

"F_1=2500\:\rm N"

"F_2=450\:\rm N"



The net force


"F=F_1-F_2=2500-450=2050\:\rm N\quad downward"